# On Biologically Plausible Learning in Continuous Time

### v1

@@ Abstract
Biological learning unfolds continuously in time, yet most algorithmic models rely on discrete updates and separate inference and learning phases. We study a continuous-time neural model that unifies several biologically plausible learning algorithms and removes the need for phase separation. Rules including stochastic gradient descent (SGD), feedback alignment (FA), direct feedback alignment (DFA), and Kolen–Pollack (KP) emerge naturally as limiting cases of the dynamics. Simulations show that these continuous-time networks stably learn at biological timescales, even under temporal mismatches and integration noise. Through analysis and simulation, we show that learning depends on temporal overlap: a synapse updates correctly only when its input and the corresponding error signal coincide in time. When inputs are held constant, learning strength declines linearly as the delay between input and error approaches the stimulus duration, explaining observed robustness and failure across network depths. Critically, robust learning requires the synaptic plasticity timescale to exceed the stimulus duration by one to two orders of magnitude. For typical cortical stimuli (tens of milliseconds), this places the functional plasticity window in the few-second range, a testable prediction that identifies seconds-scale eligibility traces as necessary for error-driven learning in biological circuits.
@@ 1 Introduction
Understanding how biological circuits learn has long been a central challenge at the interface of neuroscience and machine learning. Many algorithmic proposals for biologically plausible learning have been put forward, including feedback alignment (FA; Lillicrap et al. (2016)), direct feedback alignment (DFA; Nøkland (2016)), and Kolen–Pollack / weight mirrors (KP; Kolen and Pollack (1994); Akrout et al. (2019)). These methods attempt to relax the strong requirements of backpropagation, such as exact weight transport, while preserving the ability to optimize deep networks. Such rules are attractive both as potential models of cortical learning and as candidates for training in neuromorphic or analog hardware.
Two key aspects of real biological systems, however, are conspicuously absent from most of these formulations: learning in biology unfolds continuously in time, and does not rely on separate learning and inference phases. Realistic neurons operate with finite conduction and integration times, so neither inference nor plasticity can be assumed to occur instantaneously or in synchronized steps (Kandel et al., 2000). By contrast, existing algorithmic models are cast in discrete steps that alternate between inference and learning phases, effectively assuming that forward and backward signals are globally and instantaneously synchronized (Whittington and Bogacz, 2019). Therefore, the existing biologically plausible algorithms fail to answer a basic question: can learning be achieved in continuous time, without artificial phase separation, and at biologically realistic timescales?
In this work, we address this gap by formulating a continuous-time neural model in which both neural states and synaptic weights evolve according to coupled first-order differential equations. Each neuron receives a feedforward drive and a modulatory error drive, and updates its synapses through a local two-signal rule. Crucially, inference and learning occur simultaneously: there is no global clock and no need for phase separation as all state variables evolve under the same ODE. The model is governed by distinct time constants for fast signal propagation, intermediate plasticity, and slow synaptic decay, an ordering that mirrors the hierarchy of timescales in real neural tissue.
Our contributions are threefold:
@@
    Continuous-time formulation. We introduce a dynamical-systems model where inference and learning unfold simultaneously, without phase separation. Classical rules and biologically plausible algorithms such as SGD, FA, DFA, and KP emerge as limiting cases under timescale separation (Section 3).
@@
    Temporal learning analysis. We analyze the effect of timing mismatches and show that learning is determined by the temporal overlap between input and error. This predicts failure when delay approaches the sample duration, depth fragility from accumulated lags, and robustness patterns observed in simulation (Sections 45).
@@
    Biological timescales. We demonstrate effective learning at biologically realistic timescales and derive a quantitative constraint: plasticity windows must outlast the stimulus by an order of magnitude, placing the functional eligibility regime in the few-second range (Section 5.3).
Together, these results support the biological feasibility of rules like KP, FA, and DFA when cast in continuous time. They show that plasticity arises from temporal overlap between input and error, explaining the robustness of these algorithms and providing a physically grounded path toward continuous-time learning in both biological and artificial systems.
@@ 2 Related Work
Related families of algorithms, such as contrastive Hebbian learning (Xie and Seung, 2003), equilibrium propagation (Scellier and Bengio, 2017), and other energy-based formulations (Hopfield, 1982; Bengio and Fischer, 2015), also instantiate continuous-time learning rules. These approaches rely on symmetric connectivity or two-phase equilibrium dynamics, which differ from the heterosynaptic two-signal rules that are our focus here. While our model shares the spirit of framing learning as a dynamical process, we restrict our focus in this work to error propagation style rules (FA, DFA, KP) and their continuous-time realizations.
Both neural ODEs (Chen et al., 2018) and our neural differential-equation model cast network computation as a continuous-time dynamical system, replacing discrete layers/updates with ODE flows over time. In standard neural ODEs, parameters are fixed during the forward solve, and gradients are typically recovered by integrating an adjoint ODE backward in time (Chen et al., 2018) or by differentiating through the solver (Baydin et al., 2018). By contrast, our model couples learning and inference in one forward-in-time system: both neural states and parameters evolve by ODEs, with weights updated online via locally computed and propagated error terms.
Concurrently, “self-assembling” heterosynaptic circuits specifying four plastic synapses linking forward and feedback streams have been shown to match backpropagation on standard benchmarks (Liao et al., 2024). A key insight from this line of work is that error-propagation rules such as feedback-style algorithms (FA, DFA, KP) can emerge without requiring explicit weight symmetry, and only using local rules. This further motivates our focus on these algorithms as being attractive candidates for biologically plausible and hardware-realizable learning.
Biological relevance of feedback-type algorithms has been argued in various works. For example, Lillicrap et al. (2020) argued how FA may be realizable in the cerebellum, where the Purkinje cells are the trainable neurons, and the one-to-one climbing fibres carry in the feedback error signals. Koplow et al. (2025) computationally showed how these algorithms are consistent with the widely observed Hebbian and anti-Hebbian plasiticities in the brain. However, it is not yet clear whether these algorithm can work in the continuous-time, and if so, whether they work at biologically reasonable timescales. This is a key question we address.
3 A Continuous-Time Neural Network Model
@@
Figure 1 illustrates the continuous-time heterosynaptic neuronal systems we study. The forward pathway, parameterized by $\mathbf{w}$ and with activation $\sigma$, maps inputs $\mathbf{x}$ to neuronal output $z$. The error pathway, parameterized by $\mathbf{v}$, receives an error signal $\mathbf{e}$ and provides a modulatory influence on learning. Learning is heterosynaptic: the forward weights $\mathbf{w}$ update in proportion to the modulatory drive $(\mathbf{v}^{\top}\mathbf{e})$, while the error weights $\mathbf{v}$ update in proportion to the forward drive $(\mathbf{w}^{\top}\mathbf{x})$. Let layer $l$ have width $d_{l}$. Stacking the per-neuron updates, we can use the following notation: $\mathbf{z}_{l-1}(t)\in\mathbb{R}^{d_{l-1}}$, $\mathbf{z}_{l}(t)\in\mathbb{R}^{d_{l}}$, $W_{l}\in\mathbb{R}^{d_{l}\times d_{l-1}}$, and error source $e_{l}^{\text{src}}(t)$. Define
@@
$\dot{\mathbf{z}}_{l}=\frac{-\mathbf{z}_{l}+\sigma_{l}(W_{l}^{\top}\mathbf{z}_{l-1})}{\tau_{\mathrm{prop}}},\quad\dot{W}_{l}=-\frac{W_{l}}{\tau_{\mathrm{dec}}^{W}}+\frac{\mathbf{z}_{l-1}\,(V_{l}^{\top}\bm{\epsilon}_{l})^{\top}}{\tau_{\mathrm{plas}}^{W}},\quad\dot{V}_{l}=-\frac{V_{l}}{\tau_{\mathrm{dec}}^{V}}+\frac{(W_{l}^{\top}\mathbf{z}_{l-1})\,\bm{\epsilon}_{l}^{\top}}{\tau_{\mathrm{plas}}^{V}}.$ (1)
@@
Here $\tau_{\mathrm{prop}}$ is the neuronal propagation time constant, setting how quickly $\mathbf{z}_{l}$ relaxes to its driven input. The constants $\tau_{\mathrm{plas}}^{W},\tau_{\mathrm{plas}}^{V}$ control how rapidly synaptic plasticity occurs when presynaptic and modulatory drives coincide. The constants $\tau_{\mathrm{dec}}^{W},\tau_{\mathrm{dec}}^{V}$ govern passive weight decay, setting the forgetting timescale. For the biological interpretation of these timescales, see Section 5.3. These equations arise by stacking per-neuron heterosynaptic updates; see Appendix A for the per-neuron update rules.
We emphasize that these equations describe rate-based dynamics, not spiking dynamics: the neuronal state $z(t)$ evolves continuously without threshold-triggered resets or discrete spikes (Hopfield, 1984; Wilson and Cowan, 1972). This abstraction allows us to focus on continuous-time error-propagation rules and makes the analysis tractable. While the lack of spikes is a limitation that excludes important timing effects and spike-based coding strategies, it can still serve as a reasonable model for certain classes of cortical neurons whose average firing rates are well-approximated by continuous variables (Wilson and Cowan, 1972). Moreover, rate-based neurons align naturally with the error-propagation algorithms we study (FA, DFA, KP), which are themselves formulated in terms of continuous activations rather than discrete events.
@@
These dynamics can be understood as a continuous-time realization of the heterosynaptic two-signal principle. Under separation of timescales, we may treat neuronal activity as equilibrated relative to the slower synaptic updates. In this quasi-stationary regime, the fast variables satisfy $\mathbf{z}_{l}=\sigma(W_{l}^{\top}\mathbf{z}_{l-1})$, and the weight dynamics reduce to the heterosynaptic updates analyzed in Ziyin et al. (2025), where local stability guarantees each neuron implements a gradient step and global consistency ensures alignment across layers. At full stationarity ($\dot{\mathbf{z}}=\dot{W}=\dot{V}=0$) and with equal size, the fixed-point equations imply $W_{l}\propto\mathbf{z}_{l-1}(V_{l}^{\top}\bm{\epsilon}_{l})^{\top}$ and $V_{l}\propto(W_{l}^{\top}\mathbf{z}_{l-1})\bm{\epsilon}_{l}^{\top}$, so that $W_{l}$ converges to a configuration consistent with $V_{l}^{\top}$. In this limit, the effective update to $W_{l}$ is the outer product of the input with the backpropagated error at that layer—recovering the canonical backpropagation rule as the stationary solution of the dynamics.
The center and right panels of Figure 1 show how these neurons connect into multilayer networks. Under timescale separation $\tau_{\mathrm{prop}}{}\ll\tau_{\mathrm{plas}}\ll\tau_{\mathrm{dec}}$, the activation states equilibrate quickly, and weights update quasi-statically. Outputs track their steady state $\mathbf{z}_{l}\approx\sigma_{l}(W_{l}^{\top}\mathbf{z}_{l-1})$ within a sample window. Averaging the weight ODE over one input data presentation gives a weight update $\Delta W_{l}\propto\mathbb{E}_{\text{window}}\left[\mathbf{z}_{l-1}(V_{l}^{\top}\bm{\epsilon}_{l})^{\top}\right]$. The effective weight update rule corresponding to different learning algorithms depends only on the definitions of the feedback weights $V_{l}$ and error drive $\bm{\epsilon}_{l}$. Given the error $\mathbf{e}_{L}$ as the gradient of the loss function with respect to the output layer:
SGD. Under the constraint $V_{l}=W_{l}^{\top}$, the error signals are equal to the true backpropagated error, $\bm{\epsilon}_{l}=\mathbf{e}_{l}$. Then, $\Delta W_{l}\propto\mathbb{E}[\mathbf{z}_{l-1}\mathbf{e}_{l}^{\top}]$, which is exactly the gradient of the loss.
Feedback Alignment (FA). Set $V_{l}$ to fixed random matrices that project the error at each layer backwards to the previous layer. Then, $\Delta W_{l}\propto\mathbb{E}[\mathbf{z}_{l-1}(V_{l}^{\top}\bm{\epsilon}_{l})^{\top}]$, which is the standard FA update. Gradient alignment arises empirically (Lillicrap et al., 2016).
Direct Feedback Alignment (DFA). Set $V_{l}$ to fixed random matrices. Broadcast the output-layer error $\mathbf{e}_{L}$ to all hidden layers, such that $\bm{\epsilon}_{l}=\mathbf{e}_{L}$. Then, $\Delta W_{l}\propto\mathbb{E}[\mathbf{z}_{l-1}(V_{l}^{\top}\mathbf{e}_{L})^{\top}]$, which is the DFA update. Each layer learns from a direct projection of the global error (Nøkland, 2016).
Kolen-Pollack (KP) / Weight-Mirror Methods. Allow $V_{l}$ to evolve under the full plasticity rules in equation 1. $V_{l}$ tracks the correlation between the forward drive $W_{l}^{\top}\mathbf{z}_{l-1}$ and the error $\bm{\epsilon}_{l}$. In expectation, this update combined with weight decay pushes $V_{l}$ towards $W_{l}^{\top}$. Thus KP can be viewed as a dynamical mechanism that drives $V_{l}\to W_{l}^{\top}$, in contrast to SGD which assumes this equality is enforced from the start (Akrout et al., 2019; Kolen and Pollack, 1994).
@@
4 Timing Robustness
@@
$\Delta W_{l}\;\propto\;\int_{0}^{T}\mathbf{z}_{l-1}(t)\,\big(V_{l}^{\top}\bm{\epsilon}_{l}(t)\big)^{\top}\,k_{\tau_{\mathrm{plas}}}(t)\;dt,\qquad k_{\tau_{\mathrm{plas}}}(t)=\exp\!\big(-(T-t)/\tau_{\mathrm{plas}}\big),$ (2)
@@
$\Delta(W_{l})_{ij}\;\propto\;\int_{0}^{T}(\mathbf{z}_{l-1}(t))_{i}\;\big((V_{l})_{:,j}^{\top}\bm{\epsilon}_{l}(t)\big)\;k_{\tau_{\mathrm{plas}}}(t)\;dt.$ (3)
@@
$\mathbb{E}[\Delta(W_{l})_{ij}]$ $\propto\int_{t_{0}}^{t_{1}}\exp\!\big((t-T)/\tau_{\mathrm{plas}}\big)\,dt$
@@
$\mathbb{E}[\Delta(W_{l})_{ij}]\;\propto\;(T-|\Delta|)_{+}.$ (5)
@@
so that $K_{C}(\Delta)+K_{I}(\Delta)=K_{T}:=\int_{0}^{T}k(t)\,dt$. Thus overlap and mismatch trade off under a fixed budget. In the flat-kernel case $k\equiv 1$, this reduces to L+(TL)=TL+(T-L)=T, yielding $\mathbb{E}[\Delta(W_{l})_{ij}]\propto L=(T-|\Delta|)_{+}$, i.e. accuracy improves with either larger $T$ or smaller $|\Delta|$.
@@
5 Experiments

### v2

# Does Feedback Alignment Work at Biological Timescales?

@@ Abstract
Feedback alignment and related weight-transport–free algorithms are often proposed as biologically plausible alternatives to backpropagation, yet they are typically formulated in discrete phases with implicitly synchronized forward and error signals. We develop a continuous-time model of feedback-alignment–type learning in which neural activities and synaptic weights evolve together under coupled first-order dynamics with distinct propagation, plasticity, and decay time constants. We show that learning is governed by the temporal overlap between presynaptic drive and a locally projected error signal, providing an analytic explanation for robustness to moderate timing mismatch and for failure when mismatch eliminates overlap. Our results show that in order for feedback-alignment–type algorithms to work at biological timescales, they must obey the same temporal overlap principle that applies to other biological processes like eligibility traces.
@@ 1 Introduction
Understanding how biological circuits learn has long been a central challenge at the interface of neuroscience and machine learning. Among the many proposals for biologically plausible learning, a prominent class consists of feedback-alignment–style [^1] In this work, we will use the term “feedback alignment” to refer to the class of feedback-alignment–style algorithms., weight-transport–free algorithms such as feedback alignment (FA; Lillicrap et al. (2016)), direct feedback alignment (DFA; Nøkland (2016)), and Kolen–Pollack / weight mirrors (KP; Kolen and Pollack (1994); Akrout et al. (2019)). These algorithms aim to relax the strong requirements of backpropagation, such as exact weight transport, while preserving the ability to optimize deep networks. These algorithms are attractive both as potential models of cortical learning and as approaches for training in neuromorphic or analog hardware.
There are two primary difficulties in connecting FA–type algorithms to biology: learning in biology unfolds continuously in time, and does not rely on separate learning and inference phases. Realistic neurons operate with finite conduction and integration times, so neither inference, error propagation, nor plasticity can be assumed to occur instantaneously or in synchronized steps (Kandel et al., 2000). Unlike digital learning algorithms, biological systems do not alternate between distinct inference and learning phases. By contrast, feedback alignment is cast in discrete steps that alternate between inference and learning phases, effectively assuming that forward and backward signals are globally and instantaneously synchronized (Lillicrap et al., 2016; Nøkland, 2016; Akrout et al., 2019; Scellier and Bengio, 2017). This leaves a basic question unaddressed: Do FA/KP/DFA–style learning rules still work when implemented as continuous-time processes with biologically realistic propagation and plasticity timescales?
In this work, we address this gap directly by constructing a continuous-time model for feedback-alignment–style algorithms. In this model, both neural states and synaptic weights evolve according to coupled first-order differential equations. Each neuron receives a feedforward drive and a modulatory error drive, and updates its synapses through a local two-signal rule. Crucially, inference and learning occur simultaneously: there is no global phase separation, as all state variables evolve under the same ODE. The model is governed by distinct time constants for fast signal propagation, intermediate potentiation, and slow synaptic decay, an ordering that mirrors the hierarchy of timescales in real neural tissue. Notably, feedback alignment algorithms are inherently non-spiking, and so our continuous-time model also does not attempt to model spikes.
Our model reveals three main insights about feedback alignment in continuous time:
@@
    Continuous-time realization. Feedback alignment works in continuous time: inference and learning can proceed simultaneously under coupled differential equations without explicit forward/backward phases.
@@
    Biologically realistic regimes. Effective learning occurs when propagation, plasticity, and decay constants satisfy the hierarchy of timescales $\tau_{\mathrm{prop}}\ll\tau_{\mathrm{plas}}\ll\tau_{\mathrm{dec}}$, and align with biologically observed ranges.
@@
    Overlap as the operative principle. Learning is governed by temporal overlap between presynaptic drive and error signals; when delays eliminate overlap, updates become biased and performance collapses.
@@ 2 Related Work
There has been significant interest in biologically plausible models of learning. Our focus here is on feedback-alignment–type algorithms, but many related frameworks have been explored. Continuous-time learning rules arise in contrastive Hebbian learning (Xie and Seung, 2003), equilibrium propagation (Scellier and Bengio, 2017), and broader energy-based formulations (Hopfield, 1982; Bengio and Fischer, 2015). Latent-equilibrium networks and related models derive neural and synaptic dynamics as gradient flows, enabling approximations to backpropagation (and backpropagation through time) through slow continuous dynamics (Haider et al., 2021; Ellenberger et al., 2024). While these approaches take other perspectives of learning as dynamical evolution, we concentrate on error-propagation rules—feedback alignment (FA), direct feedback alignment (DFA), and Kolen–Pollack (KP)—and their continuous-time realizations.
Biologically motivated approximations to backpropagation have also been surveyed extensively (Whittington and Bogacz, 2019), including predictive-coding and dendritic-error models (Whittington and Bogacz, 2017; Sacramento et al., 2018). These frameworks typically express neuronal dynamics in continuous time and define learning via locally computed error signals. Our work complements this literature by focusing specifically on weight-transport–free error-propagation algorithms and analyzing their formulation as coupled neural and synaptic ODEs in layered feedforward networks.
Continuous-time neural computation has also been studied in neural ODEs (Chen et al., 2018), where network computation is expressed as an ODE over hidden states and gradients are obtained via adjoint methods or solver differentiation (Baydin et al., 2018). In contrast, we consider a single forward-in-time dynamical system in which both neural states and parameters evolve continuously, with weights updated online through locally propagated error signals.
Recent work on heterosynaptic circuits has demonstrated that networks with local plasticity rules linking forward and feedback pathways can approximate backpropagation on standard benchmarks (Liao et al., 2024). These results further motivate feedback-style algorithms (FA, DFA, KP) as promising candidates for biologically grounded and hardware-compatible learning.
Finally, the biological plausibility of feedback-alignment–type algorithms has been explored in several contexts. For example, Lillicrap et al. (2020) argued how FA may be realizable in the cerebellum, and Koplow et al. (2025) showed that such rules align with observed Hebbian and anti-Hebbian plasticity motifs. However, their behavior in fully continuous-time settings and at biologically realistic timescales remains insufficiently understood. Addressing this question is a central goal of the present work.
3 A Continuous-Time Model of Feedback Alignment
@@
The FA–type algorithms feature a forward pathway that performs inference and a backward pathway that transmits error signals. While the original FA algorithm requires a fixed and randomly connected feedback pathway whose topology matches the feedforward, various variants have made the backward pathway much more flexible. For this reason, we study a broad and flexible set of connectivity topologies (see some discussions of this point in Ziyin et al. (2025)). Figure 1 illustrates the continuous-time heterosynaptic neuronal systems we study. The forward pathway, parameterized by $\mathbf{w}$ and with activation $\sigma$, maps inputs $\mathbf{x}$ to neuronal output $z$. The error pathway, parameterized by $\mathbf{v}$, receives an error signal $\bm{\epsilon}$ and provides a modulatory influence on learning. Learning is heterosynaptic: the forward weights $\mathbf{w}$ update in proportion to the modulatory drive $(\mathbf{v}^{\top}\bm{\epsilon})$, while the error weights $\mathbf{v}$ update in proportion to the forward drive $(\mathbf{w}^{\top}\mathbf{x})$. Let layer $l$ have width $d_{l}$. See Appendix A for the per-neuron update rules. Stacking the per-neuron variables into matrices, we can use the following notation: $\mathbf{z}_{l-1}(t)\in\mathbb{R}^{d_{l-1}}$, $\mathbf{z}_{l}(t)\in\mathbb{R}^{d_{l}}$, $W_{l}\in\mathbb{R}^{d_{l}\times d_{l-1}}$, and error source $\bm{\epsilon}_{l}(t)$ to define
@@
$\dot{\mathbf{z}}_{l}=\frac{-\mathbf{z}_{l}+\sigma_{l}(W_{l}\mathbf{z}_{l-1})}{\tau_{\mathrm{prop}}},\quad\dot{W}_{l}=-\frac{W_{l}}{\tau_{\mathrm{dec}}^{W}}+\frac{(V_{l}\bm{\epsilon}_{l})\mathbf{z}_{l-1}^{\top}}{\tau_{\mathrm{plas}}^{W}},\quad\dot{V}_{l}=-\frac{V_{l}}{\tau_{\mathrm{dec}}^{V}}+\frac{(W_{l}\mathbf{z}_{l-1})\,\bm{\epsilon}_{l}^{\top}}{\tau_{\mathrm{plas}}^{V}}.$ (1)
@@
Here $\tau_{\mathrm{prop}}$ is the neuronal propagation time constant, setting how quickly $\mathbf{z}_{l}$ relaxes to its driven input. The constants $\tau_{\mathrm{plas}}^{W},\tau_{\mathrm{plas}}^{V}$ control how rapidly synaptic plasticity occurs when presynaptic and modulatory drives coincide. The constants $\tau_{\mathrm{dec}}^{W},\tau_{\mathrm{dec}}^{V}$ govern passive weight decay, setting the forgetting timescale. The center and right panels of Figure 1 show how these neurons connect into multilayer networks. Under timescale separation $\tau_{\mathrm{prop}}{}\ll\tau_{\mathrm{plas}}\ll\tau_{\mathrm{dec}}$, the activation states equilibrate quickly, and weights update quasi-statically. Outputs track their steady state $\mathbf{z}_{l}\approx\sigma_{l}(W_{l}\mathbf{z}_{l-1})$ within a sample window.
@@
These dynamics can be understood as a continuous-time realization of the heterosynaptic two-signal principle. Under separation of timescales, we may treat neuronal activity as equilibrated relative to the slower synaptic updates. In this quasi-stationary regime, the fast variables satisfy $\mathbf{z}_{l}=\sigma(W_{l}\mathbf{z}_{l-1})$, and the weight dynamics reduce to the standard feedforward processing. At full stationarity ($\dot{\mathbf{z}}=\dot{W}=\dot{V}=0$) and with equal size, the fixed-point equations imply $W_{l}\propto(V_{l}\bm{\epsilon}_{l})\mathbf{z}_{l-1}^{\top}$ and $V_{l}\propto(W_{l}\mathbf{z}_{l-1})\,\bm{\epsilon}_{l}^{\top}$, so that $W_{l}$ converges to a configuration consistent with $V_{l}^{\top}$. In this limit, the effective update to $W_{l}$ is the outer product of the input with the backpropagated error at that layer, recovering a backprop-like outer-product update. The effective weight update rule corresponding to different learning algorithms depends only on the definitions of the feedback weights $V_{l}$ and error drive $\bm{\epsilon}_{l}$. In Appendix B we discuss the discrete-time limits of these dynamics.
@@
4 Continuous-time feedback alignment requires temporal overlap
@@
$\Delta W_{l}\;\propto\;\int_{0}^{T}\big(V_{l}\bm{\epsilon}_{l}(t)\big)\,\mathbf{z}_{l-1}(t)^{\top}\,k_{\tau_{\mathrm{plas}}}(t)\;dt,\qquad k_{\tau_{\mathrm{plas}}}(t)=\exp\!\big(-(T-t)/\tau_{\mathrm{plas}}\big),$ (2)
@@
$\Delta(W_{l})_{ji}\;\propto\;\int_{0}^{T}(\mathbf{z}_{l-1}(t))_{i}\;\big((V_{l})_{j,:}\,\bm{\epsilon}_{l}(t)\big)\;k_{\tau_{\mathrm{plas}}}(t)\;dt.$ (3)
@@
$\mathbb{E}[\Delta(W_{l})_{ji}]$ $\propto\int_{t_{0}}^{t_{1}}\exp\!\big((t-T)/\tau_{\mathrm{plas}}\big)\,dt$
@@
$\mathbb{E}[\Delta(W_{l})_{ji}]\;\propto\;(T-|\Delta|)_{+}.$ (5)
@@
so that $K_{C}(\Delta)+K_{I}(\Delta)=K_{T}:=\int_{0}^{T}k(t)\,dt$. Thus overlap and mismatch trade off under a fixed budget. In the flat-kernel case $k\equiv 1$, this reduces to L+(TL)=TL+(T-L)=T, yielding $\mathbb{E}[\Delta(W_{l})_{ji}]\propto L=(T-|\Delta|)_{+}$, i.e. accuracy improves with either larger $T$ or smaller $|\Delta|$.
@@
5 Feedback alignment works at biological timescales